Can this be solved? I think the question is worded wrong or is a misprint, please help.
An isosceles triangle with sides AB and BC congruent has the following lengths:
AB= x+7, AC = 3x-5
Solve for x and the length of all three sides.

Answer by Alan3354(69443)

An isosceles triangle with sides AB and BC congruent has the following lengths:
AB= x+7, AC = 3x-5
Solve for x and the length of all three sides.
------------------
BC = AB = x+7
Nothing else can be determined.